True or False:


The value of the binomial coefficient "\displaystyle{{2}\choose{10}}" is zero


1
Expert's answer
2021-09-20T16:17:39-0400

The binomial coefficient for non negative integers is given by:

"\begin{pmatrix}
 n \\
 k 
\end{pmatrix}=\begin{cases} \frac{n!}{k!(n-k)!}\quad \text{for} ~0 \leq k <n
\\ 0 \quad \quad \quad \quad\text{otherwise}\end{cases}"

Note that from the given binomial coefficient that "n=2" and "k=10"

Since "k" is not less than "n" ; therefore the value of the given binomial coefficient is equal to zero.

The given statement is true.